a body of mass 10 kg is acted upon 2 forces 32N and 3t2 . find the velocity of body after 5 sec. if its initial velocity is 10m/s

Grade:11

2 Answers

Viraj C Bukitagar
8 years ago
you know that sum of all net force
Viraj C Bukitagar
8 years ago
you know that sum of all forces is equal to mass*acceleration      therefore M*a=32+3t^2
 
you know that a=dv/dt so replace it 
the expression would be 
M*dv=(32+3t^2)dt
 
now integragte by putting LHS limits as 10 to v and RHS limits as 0 to 5 
so the answer would be 38.5m/s